find the slope of any line parrallel to the line through points (18,4)and 3,5)
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

slope=(5-4)/(3-18)
slope=1/-15=-1/15 for this slope.
Any parallel line will also have a slope=-1/15.

Answer by stanbon(57361)

find the slope of any line parrallel to the line through points (18,4)and 3,5)
------------
slope = (5-4)/(3-18) = -1/15
==========================
Cheers,
Stan H.